OpenStudy (anonymous):

2x - 5y = 3 x - 3y = 1

OpenStudy (alexandervonhumboldt2):

so

OpenStudy (alexandervonhumboldt2):

x=1+3y

OpenStudy (alexandervonhumboldt2):

2*(1+3y)-5y=3

OpenStudy (alexandervonhumboldt2):

2+6y-5y=3

OpenStudy (alexandervonhumboldt2):

y=1

OpenStudy (alexandervonhumboldt2):

x=4

OpenStudy (alexandervonhumboldt2):

(4, 1)

OpenStudy (anonymous):

ty

Nnesha (nnesha):

@AlexandervonHumboldt2 don't you know direct answer not allowed ???

OpenStudy (alexandervonhumboldt2):

ishowed steps stop pretending

Nnesha (nnesha):

i'm not pretending i guess you know what are rules